OpenStudy (anonymous):

3y^2 - 13y-10 - Factoring Trinomials where a does not equal 1

OpenStudy (saifoo.khan):

\[(y-5) (3 y+2) \]

OpenStudy (anonymous):

there is a method; do you want to learn

OpenStudy (anonymous):

there and is trial and error

OpenStudy (anonymous):

Yes please. I would like to learn the method.

OpenStudy (anonymous):

if there are factors of ac that sum to b, then the trinomial is factorable\[ax^2+bx+c\]in this case ac=3*(-10)=130 and the factors are 1; 30 2; 15 3; 10 5; 6 where each pair of factors must have differing signs. (continued)

OpenStudy (anonymous):

the pair -15*2=-30 and -15+2=-13 so we can rewrite the -13y as -15y+2y to get\[3y^3-15y+2y-10\] (continued)

OpenStudy (anonymous):

now there is a method of factoring called "factoring by grouping" where you factor out the GCF of the first two terms and the GCF out of the last two terms:\[3y(y-5)+2(y-5)\](continued) NOTE: I accidentally wrote y^3 instead of y^2 for the first term in the previous post!!

OpenStudy (anonymous):

now id you think of this expression as having two terms each with a GCF of the binomial y-5, you can factor it out, leaving the 3y and the +2:\[(y-5)(3y+2)\]and this is the factored form

OpenStudy (anonymous):

same as saifoo got using trial and error most likely which just takes some practice

OpenStudy (anonymous):

Thank you for your answer. This will help me out greatly.

OpenStudy (anonymous):

you want to do another; i've got one ready to do... with less explanation

OpenStudy (anonymous):

Sure. By the way, i am new to this.

OpenStudy (anonymous):

Factor \[6x^2+x-12\]

OpenStudy (anonymous):

ac=6(-12)=-72 -8*9=-72 -8+9=1=b, thus\[6x^2-8x+9x-12\]factor out the GCF from each group\[=2x(3x-4)+3(3x-4)\]factor out the binomial\[=(3x-4)(2x+3)\]done